A gasoline tank is 1/4 full. After adding 10 gallons of gas the gauge now indicates, the tank is 2/3 full. What is the capacity of the tank.
I get that 2/3 = 10 gallons and that by going from 1/4 to 2/3 means that that 10 gallons was close to half of the tanks capacity. Please help me see what I am doing wrong?

Answer by stanbon(75887)

A gasoline tank is 1/4 full. After adding 10 gallons of gas the gauge now indicates, the tank is 2/3 full. What is the capacity of the tank.
-----
Equation:
Let "x" be the capacity of the tank.
----
(1/4)x + 10 = (2/3)x
-----
Multiply thru by 12 to get:
3x + 120 = 8x
5x = 120
x = 24 gallons (capacity of the tank)
==========================================
Cheers,
Stan H.
=================